OpenStudy (anonymous):

Where is the vertex of the parabola? y = x^2 + 2x – 3 A. Above the x-axis B. Below the x-axis C. On the x-axis D. On the y-axis

OpenStudy (anonymous):

This is the same thing I just did I think, so use the graphs from the previous question's correct answer.

OpenStudy (anonymous):

Previous one was y = 2x^2 + 5x - 3

OpenStudy (anonymous):

Oh, well then it's the same concept.

OpenStudy (anonymous):

In other words, find the point where the slope is zero.

OpenStudy (anonymous):

You can simplify your function if you'd like and locate the two coordinates where the parabola touches on the x-intercept.

OpenStudy (anonymous):

For example, y = x^2 + 2x - 3 can be transformed into y = (x - 1) ( x + 3) Therefore, the parabola goes through 1 and -3 on the x-intercept.